A fair spinner has 9 equal sections: 3 red, 2 blue and 4 green.
It is spun twice.
What is the probability of getting red both times?
1
Expert's answer
2020-05-05T19:57:36-0400

"Let\; p(red)=\frac{3}{9}, p(blue)=\frac{2}{9}, p(green)=\frac{4}{9}\\
\text{Then, the probability of getting red both times is}\\
p(red,red)=\frac{3}{9} \times \frac{3}{9}=\frac{9}{81}=\frac{1}{9}"